# DisTaC: Conditioning Task Vectors via Distillation for Robust Model Merging

### v1

@@ 3.2 Low-Confidence Source Models
We now show that low confidence constitutes a second, equally damaging failure mode. Paradoxically, models that are well calibrated can be fragile from the perspective of model merging; conversely, we argue that the more over-confident a source model is, the more robust it becomes to merging.
@@
Figure 1(a) (yellow bars) shows how this reduced confidence affects merging. Across all algorithms, the normalized accuracy drops markedly—by as much as 24% compared to the no-smoothing baseline. This degradation exceeds that caused by norm discrepancies in the previous section, underscoring how harmful low-confidence source models can be. In short, routine training choices that alter confidence (e.g. label smoothing, Mixup, focal loss) can induce large swings in post-merge performance. These phenomena can also be supported from a theoretical perspective. (Appendix A)
@@ 4 Knowledge Distillation for Task vector Conditioning
6:  Sample mini-batch $\mathcal{B}\subset\tilde{\mathcal{D}}^{u}_{t}$
@@
8:  for all $x\in\mathcal{B}$ do
9:   $\bm{z}_{\text{tcr}}\leftarrow f(x;\,\bm{\theta}_{\text{pre}}+\bm{\tau}_{t})$
10:   $\bm{z}_{\text{stu}}\leftarrow f(x;\,\bm{\theta})$
@@
15:  $L\leftarrow\frac{L}{|\mathcal{B}|}+\beta\lVert\bm{\theta}-\bm{\theta}_{0}\rVert_{2}^{2}$
16:  $\bm{\theta}\leftarrow\bm{\theta}-\eta\,\nabla_{\bm{\theta}}L$triangleright Gradient step
@@
Here, we propose DisTaCDistillation for Task vector Conditioning—a KD–based pre-conditioning method that eliminates the harmful effects of individual task vectors during model merging, as identified in Section 3.
@@ 4.1 Task Vector Norm Conditioning
Although one might instead fine-tune $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}_{t}$ with labeled examples, obtaining a sufficiently large supervised corpus at merge time is typically impractical. By contrast, access to unlabeled data is commonly assumed during model merging (Yang et al., 2024b; Yan et al., 2025; Yoshida et al., 2025), and KD imposes only mild additional requirements.
@@ 4.2 Source Model Confidence Conditioning
One may worry that the over-confidence harms model reliability in practice. However, standard post-hoc calibration methods (e.g. temperature scaling) can mitigate over-confidence, whereas merging with an under-confident model leads to large performance drops that make the merged model impractical. A detailed discussion appears in Section 6.2.
@@ 5.1 Setup
We conducted experiments in a multitask setting following Ilharco et al. (2023). Specifically, we adopted eight vision tasks—Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SUN397 (Xiao et al., 2016), and SVHN (Netzer et al., 2011). Our models applied ViT-B-32 and ViT-L-14 to CLIP. We evaluated post-merge performance using absolute accuracy and normalized accuracy under the two aforementioned failure modes: the case with diverged task vector norms (Norm Mismatch) and the case with low-confidence source models (Low Confidence). The detailed settings for each scenario followed those described in Section 3. We adopt four merging methods as baselines: task arithmetic (Ilharco et al., 2023), Ties-Merging (TIES) (Yadav et al., 2023), Consensus Merging (Consensus TA)(Wang et al., 2024), and TSVM (Gargiulo et al., 2025). For DisTaC, knowledge distillation was run for $K=500$ steps. In the Norm Mismatch regime we assign a task–specific scaling coefficient $\kappa_{t}$ individually for each of the eight norm–disparity configurations: the task vector with the largest $\ell_{2}$-norm is rescaled so that, after scaling, its norm equals the mean norm of the remaining seven task vectors. A neutral temperature pair is then used, $(T_{\text{tcr}},T_{\text{stu}})=(10,10)$. In the Low Confidence regime we instead fix $\kappa_{t}=1$ and sharpen the student by adopting a more asymmetric temperature pair, $(T_{\text{tcr}},T_{\text{stu}})=(1,10)$. More detailed settings can be found in Appendix B.
@@ 5.2.1 Merging Performance
Table 1 summarizes the results. Absolute accuracy is displayed in a larger font, whereas normalized accuracy appears in parentheses in a smaller font. As noted in Section 3, all methods exhibit a substantial and consistent performance decline relative to the conventional configuration (Original) under both failure modes, revealing a clear vulnerability (white rows). The rows highlighted in gray show the performance obtained by first applying DisTaC for pre-conditioning and then merging. DisTaC consistently enhances merge performance, yielding gains of up to 20.8% absolute accuracy for ViT-B-32 and 18.1% for ViT-L-14. Moreover, for TSVM, which achieves the highest merge performance, DisTaC raises the accuracy under both failure modes to a level comparable with the Original configuration, indicating that the intended merge performance is robustly maintained even in challenging scenarios.
@@ 5.2.2 Efficiency of DisTaC
First, Figure 2(a) depicts the training history under the Norm Mismatch setting in Table 1. It achieves performance comparable to, or even surpassing, the teacher model’s test performance within 500 steps, while the $\ell_{2}$ regularizer of DisTaC keeps the task vector norm to roughly $\times1$ that of the initialization point, $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$, at the end of the 500 steps.
Of particular interest is that DisTaC occasionally surpasses the teacher model’s test performance. We identify two factors underlying this phenomenon. The first is the scale given by $\kappa_{t}$. In particular, we observed that reducing $\kappa_{t}$ can sometimes improve generalization performance. That is, the DisTaC initialization point already outperforms the teacher model, and we observed this in every instance in which the teacher model was exceeded. This phenomenon of the student outperforming the teacher is confirmed in (Furlanello et al., 2018), where it has been shown that a student can surpass the teacher by repeating KD between identical architectures. Furthermore, in this case, since KD is performed while keeping the student’s norm smaller than the teacher’s, it is plausible that a regularization effect similar to weight decay is being exhibited.
Next, Figure 2(b) presents the training history under the Low Confidence setting in Table 1. Within 500 steps—particularly during the first 100 steps—it achieves a substantial reduction in prediction entropy while maintaining test accuracy at a level nearly equivalent to that of the teacher model.
@@ 6 Discussion
Figure 3: Effect of scaling task vectors on test accuracy. For each of the eight vision tasks (ViT-B-32), we evaluate the model $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$ as the scaling factor $\kappa_{t}$ varies from 0.00.00.0 to 3.03.03.0. Model performance is more robust to shrinking the task vector than to stretching it, suggesting that when harmonizing task vector norms, longer vectors should be shrunk to match shorter ones.
When task vectors differ significantly in norm, a natural question arises: Should shorter vectors be stretched to match longer ones, or should longer vectors be shrunk to match the shorter ones? Our findings support the latter—we advocate shrinking the longer vectors.
There are several reasons for this. First, it is conceivable that model performance is more robust to scaling down a task vector than scaling it up. Figure 3 shows how test accuracy varies across vision tasks when applying different scaling factors $\kappa_{t}$ to the task vector, i.e., evaluating $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$ for $\kappa_{t}\in[0.0,3.0]$. Shrinking the task vector ($\kappa_{t}<1.0$) retains performance comparable to or even better than the original fine-tuned model across a broad range. In contrast, stretching beyond $\kappa_{t}=1.0$ degrades accuracy, and by $\kappa_{t}=3.0$, the model underperforms even the zero-shot baseline across all tasks. A similar trend was also observed for ViT-L-14 (see Section C.3).
As shown earlier in Figure 1(b), real-world fine-tuning pipelines often result in over 5×5$\times$5 × variation in task vector norm due to differing learning rates or training durations. In such cases, stretching small-norm vectors to match larger ones risks disrupting the pretrained model’s useful representations and is therefore undesirable.
Furthermore, Ilharco et al. (2023) observed that merging task vectors with smaller norms tends to yield better performance. A likely explanation is that smaller displacements remain within the local linear regime around $\bm{\theta}_{\text{pre}}$, where first-order approximations hold more accurately. This also aligns with the NTK perspective discussed in Ortiz-Jimenez et al. (2023); Yoshida et al. (2025), under which merging remains valid and weight disentanglement is preserved near the pretrained initialization.
@@ 6.2 Confidence Reliability in Model Merging
Figure 4: Impact of label smoothing on confidence calibration and merge performance. (a) Average reliability diagram for ViT-B-32 across eight vision tasks under different label-smoothing strengths $\alpha$. Without label smoothing ($\alpha=0$, dark purple) the model is strongly over-confident; as $\alpha$ increases to 0.010.010.01 the model becomes well-calibrated, and at $\alpha=0.1$ it turns under-confident. (b) Test normalized accuracy obtained when the corresponding source models are merged. Merge performance decreases monotonically with larger $\alpha$, revealing a clear trade-off: lower confidence comes at the cost of lower accuracy after merging.
@@
First, the calibration curves in Figure 4(a) show that a model trained without label smoothing (dark-purple line) is strongly over-confident—consistent with the well-known tendency of modern deep networks (Guo et al., 2017). As $\alpha$ increases from 0.01 to 0.1 (red $\to$ yellow), the models become well-calibrated and eventually under-confident, matching the observations of Müller et al. (2019).
Figure 4(b) then reports the normalized accuracy obtained when these source models are merged. Accuracy decreases monotonically with larger $\alpha$, revealing an inverse correlation between label-smoothing strength and merge performance.
@@
In short, current merging methods perform best when the source models are deliberately over-confident. To retain reliable confidence after merging, we therefore advocate applying post-hoc calibration—such as temperature scaling (Guo et al., 2017)—to the merged model rather than trying to calibrate the sources beforehand.
7 Related Work
7.1 Model Merging and Task Arithmetic
Research on integrating multiple neural network models by performing operations on their parameters has been widely conducted, starting with Utans (1996). These techniques enable a model to learn diverse tasks with less time and computational resources, and have become increasingly important in recent years as the number of model parameters has grown dramatically. For instance, in early approaches to model merging, models with the same architecture were fine-tuned and then merged by averaging their parameters (Wortsman et al., 2022a; Choshen et al., 2022). More sophisticated methods have since been proposed, such as Fisher Merging (Matena & Raffel, 2022), which is based on maximizing the posterior probability of the model, and RegMean (Jin et al., 2023), which minimizes the distance between output activations before and after merging. In contrast, task arithmetic (Ilharco et al., 2023) focuses on the task vector, defined as the difference in parameters between a fine-tuned model and a pre-trained model, and performs addition and subtraction of task vectors in parameter space. This approach offers the advantage of allowing flexible, localized modifications to the model and has found applications across diverse tasks (Huang et al., 2024; Su et al., 2024; Yoshikawa et al., 2025; Naganuma et al., 2025).
Recent research on task arithmetic has theoretically analyzed the simple addition of task vectors and proposed multiple methods to address its shortcomings. Approaches aimed at improving the properties of task vectors focus on the linearity in fine-tuning (Ortiz-Jimenez et al., 2023; Yoshida et al., 2025). These methods, based on the Neural Tangent Kernel (NTK) (Jacot et al., 2018), treat the model’s output as linear during fine-tuning in order to reflect vector operations in parameter space onto the model’s inputs and outputs. Meanwhile, several studies have been conducted from the perspective of mitigating interference between task vectors. TIES-Merging (Yadav et al., 2023) emphasizes the removal of redundant elements and the consideration of sign in each vector dimension. AdaMerging (Yang et al., 2024b), on the other hand, automatically adjusts merging coefficients per task and per layer to reduce task interference and enhance robustness through test-time adaptation. Wang et al. (2024) introduced a framework for pinpointing the parameters that carry information shared across tasks and, on that basis, proposed Consensus Merging, which builds task-wise masks that align more closely with inter-task consensus than the masks used in TIES-Merging. More recently, TSVM (Gargiulo et al., 2025) interprets task interference as non-orthogonality among the layer-wise singular vectors of the task vectors; by whitening those singular directions, TSVM further improves merge quality.
Despite these advances, Ilharco et al. (2023) and nearly all follow-up studies on multi-task model merging benchmark their methods under highly idealized settings, leaving real-world failure modes largely unexplored. In this work, we show that (i) discrepancies in task vector norms and (ii) low source-model confidence are key sources of interference. We introduce DisTaC as a simple pre-conditioning step that mitigates both problems before merging.
7.2 Knowledge Distillation
DisTaC addresses the limitations of existing task arithmetic methods by incorporating knowledge distillation. Knowledge distillation is a technique proposed for transferring knowledge from a teacher model to a smaller student model (Hinton et al., 2015). Although initially intended for model compression (Hinton et al., 2015; Kim et al., 2018; Sanh et al., 2020), it has also been applied in contexts such as self-distillation, where repeated distillation between models of the same architecture leads to performance improvement (Furlanello et al., 2018; Zhang et al., 2019; Zhang & Sabuncu, 2020). Among these applications, several studies have explored generating models that can handle multiple tasks by distilling knowledge from multiple teacher models (Luo et al., 2019; Hao et al., 2023). These approaches achieve distillation by mapping the parameters of multiple teacher models into a shared space for the student model. Conversely, it is also possible to distill models with different architectures individually to obtain task vectors, which can then be merged using task arithmetic (Merugu et al., 2025). DisTaC adopts the latter approach and resolves the issue of variability in the norms of task vectors by obtaining them through distillation.
Applying distillation to task arithmetic requires addressing the impact of soft targets. Numerous studies have analyzed the effects of label smoothing in the context of knowledge distillation (Müller et al., 2019; Shen et al., 2021; Chandrasegaran et al., 2022; Zheng & YANG, 2024). In this study, we demonstrate that fine-tuning with soft targets significantly affects the models obtained through mode merging, and propose a method to mitigate this effect by increasing the confidence of the student model.

### v2

@@ 3.2 Low-Confidence Source Models
We now show that low confidence constitutes a second, equally damaging failure mode. Paradoxically, models that are well calibrated can be fragile from the perspective of model merging; conversely, we argue that the more overconfident a source model is, the more robust it becomes to merging.
@@
Figure 1(a) (yellow bars) shows how this reduced confidence affects merging. In all algorithms, the normalized accuracy decreases markedly by up to 24% compared to the baseline without smoothing. This degradation exceeds that caused by norm discrepancies in the previous section, underscoring how harmful low-confidence source models can be. In short, routine training choices that alter confidence (e.g. label smoothing, Mixup, focal loss) can induce large swings in post-merge performance. These phenomena can also be supported from a theoretical perspective. (Appendix C)
@@ 4 Knowledge Distillation for Task Vector Conditioning
6:  Sample mini-batch $\mathcal{B}_{t}\subset\tilde{\mathcal{D}}^{u}_{t}$
@@
8:  for all $\bm{x}_{t}\in\mathcal{B}_{t}$ do
9:   $\bm{z}_{\text{tcr}}\leftarrow f(\bm{x}_{t};\,\bm{\theta}_{\text{pre}}+\bm{\tau}_{t})$
10:   $\bm{z}_{\text{stu}}\leftarrow f(\bm{x}_{t};\,\bm{\theta})$
@@
15:  $L\leftarrow\frac{L}{|\mathcal{B}_{t}|}+\beta\lVert\bm{\theta}-\bm{\theta}_{0}\rVert_{2}^{2}$
16:  $\bm{\theta}\leftarrow\bm{\theta}-\eta\,\nabla_{\bm{\theta}}L$ $\triangleright$ Gradient step
@@
Here, we propose Distillation for Task vector Conditioning (DisTaC) a KD–based pre-conditioning method that eliminates the harmful effects of individual task vectors during model merging, as identified in Section 3.
@@ 4.1 Task Vector Norm Conditioning
Although one might instead fine-tune $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}_{t}$ with labeled examples, obtaining a sufficiently large supervised corpus at merge time is typically impractical. By contrast, access to unlabeled data is commonly assumed during model merging (yang2024adamerging; yan2025calm; yoshida2025mastering), and KD imposes only mild additional requirements.
@@ 4.2 Source Model Confidence Conditioning
One may worry that the over-confidence harms model reliability in practice. However, standard post-hoc calibration methods (e.g. temperature scaling) can mitigate over-confidence, whereas merging with an underconfident model leads to large performance drops that make the merged model impractical. A detailed discussion appears in Section 6.2.
@@ 5 Experiment
Iso-CTS 81.0 (89.7) 90.4 (96.4) 78.1 (86.2) 90.8 (96.9) 72.5 (81.1) 80.8 (86.0)
Iso-CTS + DisTaC - - 80.3 (88.9) 90.1 (96.1) 69.0 (78.1) 86.1 (91,5)
WUDI-Merging 85.5 (93.9) 91.7 (97.7) 49.2 (52.6) 57.9 (60.8) 38.0 (40.8) 28.0 (29.2)
WUDI-Merging + DisTaC - - 84.4 (93.2) 91.4 (97.5) 73.8 (83.3) 91.6 (97.3)
@@ 5.1 Setup
We conducted experiments in a multi-task setting following ilharco2023editingmodelstaskarithmetic. Specifically, we adopted eight vision tasks: Cars (krause20133d), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), GTSRB (stallkamp2011german), MNIST (lecun1998mnist), RESISC45 (cheng2017remote), SUN397 (xiao2016sun), and SVHN (netzer2011reading). Our models applied ViT-B-32 and ViT-L-14 to CLIP. We evaluated post-merge performance using absolute accuracy and normalized accuracy under the two aforementioned failure modes: the case with diverged task vector norms (Norm Mismatch) and the case with low-confidence source models (Low Confidence). The detailed settings for each scenario followed those described in Section 3. We adopted four merging methods as baselines: task arithmetic (ilharco2023editingmodelstaskarithmetic), Ties-Merging (TIES) (yadav2023tiesmerging), Consensus Merging (Consensus TA) (wang2024localizing), EMR-Merging (huang2024emr), TSVM (gargiulo2025task), Iso-Merging (Iso-CTS) (marczak2025notaskleftbehind), and WUDI-Merging (cheng2025whoever). For DisTaC, knowledge distillation was run for $K=500$ steps. In the Norm Mismatch regime we assign a task–specific scaling coefficient $\kappa_{t}$ individually for each of the eight norm–disparity configurations: the task vector with the largest $\ell_{2}$-norm is rescaled so that, after scaling, its norm equals the mean norm of the remaining seven task vectors. A neutral temperature pair is then used, $(T_{\text{tcr}},T_{\text{stu}})=(10,10)$. In the Low Confidence regime we instead fix $\kappa_{t}=1$ and sharpen the student by adopting a more asymmetric temperature pair, $(T_{\text{tcr}},T_{\text{stu}})=(1,10)$. More detailed settings can be found in Appendix D.
@@ 5.2.1 Merging Performance
Table 1 summarizes the results. Absolute accuracy is displayed in a larger font, whereas normalized accuracy appears in parentheses in a smaller font. As noted in Section 3, all methods exhibit a substantial and consistent performance decline relative to the conventional configuration (Original) under both failure modes, revealing a clear vulnerability (white rows). The rows highlighted in gray show the performance obtained by first applying DisTaC for pre-conditioning and then merging. DisTaC consistently enhances merge performance, yielding gains of up to 35.8% absolute accuracy for ViT-B-32 and 63.6% for ViT-L-14. Moreover, for EMR-Merging, which achieves the highest merge performance, DisTaC raises the accuracy under both failure modes to a level comparable with the Original configuration in most cases, indicating that the intended merge performance is robustly maintained even in challenging scenarios.
@@ 5.2.2 Efficiency of DisTaC
First, Figure 2(a) depicts the training history under the Norm Mismatch setting in Table 1. It achieves performance comparable to, or even surpassing, the teacher model’s test performance within 500 steps, while the $\ell_{2}$ regularizer of DisTaC keeps the task vector norm to roughly $1.1\times$ that of the initialization point, $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$, at the end of the 500 steps.
Of particular interest is that DisTaC occasionally surpasses the teacher model’s test performance. We identify two factors underlying this phenomenon. The first is the scale given by $\kappa_{t}$. In particular, we observed that reducing $\kappa_{t}$ can sometimes improve generalization performance. That is, the DisTaC initialization point already outperforms the teacher model, and we observed this in every instance in which the teacher model was exceeded. This phenomenon of the student outperforming the teacher is confirmed in (furlanello2018born), where it has been shown that a student can surpass the teacher by repeating KD between identical architectures. Furthermore, in this case, since KD is performed while keeping the student’s norm smaller than the teacher’s, it is plausible that a regularization effect similar to weight decay is being exhibited.
Next, Figure 2(b) presents the training history under the Low Confidence setting in Table 1. Within 500 steps, particularly during the first 100 steps, it achieves a substantial reduction in prediction entropy while maintaining test accuracy at a level nearly equivalent to that of the teacher model.
@@ 6 Discussion
Figure 3: Effect of scaling task vectors on test accuracy. For each of the eight vision tasks (ViT-B-32), we evaluate the model $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$ as the scaling factor $\kappa_{t}$ varies from $0.0$ to $3.0$. Model performance is more robust to shrinking the task vector than to stretching it, suggesting that when harmonizing task vector norms, longer vectors should be shrunk to match shorter ones.
When task vectors differ significantly in norm, a natural question arises: Should shorter vectors be stretched to match longer ones, or should longer vectors be shrunk to match the shorter ones? Our findings support the latter; we advocate shrinking the longer vectors.
There are several reasons for this. First, it is conceivable that model performance is more robust to scaling down a task vector than scaling it up. Figure 3 shows how test accuracy varies across vision tasks when applying different scaling factors $\kappa_{t}$ to the task vector, i.e., evaluating $\bm{\theta}_{\text{pre}}+\kappa_{t}\bm{\tau}$ for $\kappa_{t}\in[0.0,3.0]$. Shrinking the task vector ($\kappa_{t}<1.0$) retains performance comparable to or even better than the original fine-tuned model across a broad range. In contrast, stretching beyond $\kappa_{t}=1.0$ degrades accuracy, and by $\kappa_{t}=3.0$, the model underperforms even the zero-shot baseline across all tasks. A similar trend was also observed for ViT-L-14 (see Section E.3).
As shown earlier in Figure 1(b), real-world fine-tuning pipelines often result in over $5\times$ variation in task vector norm due to differing learning rates or training durations. In such cases, stretching small-norm vectors to match larger ones risks disrupting the pretrained model’s useful representations and is therefore undesirable.
Furthermore, ilharco2023editingmodelstaskarithmetic observed that merging task vectors with smaller norms tends to yield better performance. A likely explanation is that smaller displacements remain within the local linear regime around $\bm{\theta}_{\text{pre}}$, where first-order approximations hold more accurately. This also aligns with the NTK perspective discussed in ortiz-jimenez2023task; yoshida2025mastering, under which merging remains valid and weight disentanglement is preserved near the pretrained initialization. Notably, Theorem 3.1 in wei2025unifying demonstrates that the performance gap between the merged model and the fine-tuned model is proportional to the product of the learning rate and the number of fine-tuning steps. This theoretical insight aligns with our claim that shrinking task vectors is preferable.
@@ 6.2 Confidence Reliability in Model Merging
Figure 4: Impact of label smoothing on confidence calibration and merge performance. (a) Average reliability diagram for ViT-B-32 across eight vision tasks under different label-smoothing strengths $\alpha$. Without label smoothing ($\alpha=0$, dark purple) the model is strongly overconfident; as $\alpha$ increases to $0.01$ the model becomes well-calibrated, and at $\alpha=0.1$ it turns underconfident. (b) Test normalized accuracy obtained when the corresponding source models are merged. Merge performance decreases monotonically with larger $\alpha$, revealing a clear trade-off: lower confidence comes at the cost of lower accuracy after merging.
@@
First, the calibration curves in Figure 4(a) show that a model trained without label smoothing (dark-purple line) is strongly overconfident, which is consistent with the well-known tendency of modern deep networks (guo2017calibration). As $\alpha$ increases from 0.01 to 0.1 (red $\to$ yellow), the models become well-calibrated and eventually underconfident, matching the observations of muller2019does. Figure 4(b) then reports the normalized accuracy obtained when these source models are merged. Accuracy decreases monotonically with larger $\alpha$, revealing an inverse correlation between label-smoothing strength and merge performance.
@@
In short, current merging methods perform best when the source models are deliberately overconfident. To retain reliable confidence after merging, we therefore advocate applying post-hoc calibration, such as temperature scaling (guo2017calibration), to the merged model rather than trying to calibrate the sources beforehand.